In a particular suburb of Melbourne, 50% of the inhabitants speak English as a first language, 20% speak Chinese, 15% Greek, 10% Polish, and 5% any other language. At the entrance of the local shopping centre, somebody is asking ten random people a number of questions.

Calculate the probability that:

• Two of these people speak Polish and two speak Greek as a first language.

• Three of these ten people speak Chinese as a first language.

• At least four of these ten people speak English as a first language.

• At most two of these ten people speak Chinese as a first language.
Answer #1

We are given here that:
P( english) = 0.5, P( chinese) = 0.2, P(greek) = 0.15, P(polish) = 0.1 and P(others) = 0.05

n = 10 is the random sample size here of questions

a) Probability that two of the 10 people speak polish, and 2 speak greek as the first language is computed here as:

= Number of ways to choose 2 people out of 10 people who speaks Polish * ( P(polish) )2 * Probability to choose 2 people out of the 8 people who speaks Greek * ( P(greek))2 * [1 - P(polish) - P(greek)]6

= (10)0.12 + (%)0.152 + (1 – 0.1 – 0.15)= 0.0505

Therefore 0.0505 is the required probability here.

b) Probability that 3 of the 10 randomly people selected speak chinese is computed here as:

= (19) 0.28(1 – 0.2)? = 0.2013

Therefore 0.2013 is the required probability here.

c) Probability that at least 4 of the 10 people speak english as a first language is computed as:

X ~ Bin(n = 10,p=0.5)

P(X > 4) = 1 - P(X = 0) - P(X = 1) - P(X = 2) - P(X = 3)

/10 P(X > 4) = 1 – 0.510 (1 + 10 +

P(X > 4) = 0.828125

Therefore 0.828125 is the required probability here.

d) Probability that at most 2 of the 10 people speak chinese is computed here as:

PY < 2) = P(Y = 0) +PY = 1) + P(Y = 2)

710 PY <2) = 0.810 + 10 * 0.2* 0.89 + | 0.2²*0.88

PY <2) = 0.6778

Therefore 0.6778 is the required probability here.